OpenStudy (anonymous):

Find the inverse of f(x) = 4x + 12

jimthompson5910 (jim_thompson5910):

f(x) = 4x + 12 y = 4x + 12 ... replace f(x) with y x = 4y + 12 ... swap x and y

jimthompson5910 (jim_thompson5910):

now solve for y to get your inverse function

OpenStudy (anonymous):

y = 4x + 12 Swapping x and y. x = 4y + 12 Now solve for y = again. 4y = x-12 4y/4 = (x-12)/4 y = x/4 - 3 Voila. There's your inverse equation. (I hope I got it right anyway.)